OpenStudy (anonymous):

Find the volume of the solid when rotated about the y-axis. y=3-x^2/9 x=3 the answer is 45/2 pi can you show me how to get that answer? thank you

OpenStudy (anonymous):

If you are rotating about the \(y\) axis, you want the function in terms of \(x\).

OpenStudy (anonymous):

Then you want to integrate with respect to \(y\).

OpenStudy (anonymous):



OpenStudy (anonymous):

integrate from x=0 to x=3

OpenStudy (anonymous):

\[ y=3-\frac{x^2}{9}\implies x^2 = 9(3-y) = 27-9y \]

OpenStudy (anonymous):

so x^2= -9y+27 and x=3